A helicopter, starting from rest, accelerates straight up from the roof of a hospital. The lifting force does work in raising the helicopter. An 760-kg helicopter rises from rest to a speed of 10.0 m/s in a time of 5.0 s. During this time it climbs to a height of 8.9 m. What is the average power generated by the lifting force?

Bicyclists in the Tour de France do enormous amounts of work during a race. For example, the average power per kilogram generated by Lance Armstrong (m = 75.0 kg) is 6.50 W per kilogram of his body mass. (a) How much work does he do during a 187-km race in which his average speed is 11.2 m/s? (b) Often, the work done is expressed in nutritional Calories rather than in joules. Express the work done in part (a) in terms of nutritional Calories, noting that 1 joule = 2.389 x 10-4 nutritional Calories.
Answer #1

classmate Date Page Ans: Pastas About Helicoptes- Taking upwards disection downward direction as Negative as positive and Masdussmate continued Date Page - Total Work done by mg + N Total Work done by f = Ak OR 66 297.3 +W: 38000 W = (38000 + 66287.9continued Hence <= 187000 S At = 16696.42865. Now let W be the total wolt dome then. P = W W = PAT At Hence- W=(487.5 X 166.9